OpenStudy (anonymous):

Find the value of x so that triangle EFFG is congruent to triangle GHE by LA.

OpenStudy (anonymous):

Plz help me understand

OpenStudy (anonymous):

@Harkirat Can you help me out :)

OpenStudy (anonymous):

let me check..

OpenStudy (anonymous):

for triangles to be congruent, the two angles have to be equal in measurement so 10x - 8 = 8x + 6 10x - 8x = 6 + 8 2x = 14 x = 14/2 x = 7 so correct answer is x = 7

OpenStudy (anonymous):

Oh thank god xD i just messed up on my math I though I was lost, THank you